# Multimodal Self-Instruct: Synthetic Abstract Image and Visual Reasoning Instruction Using Language Model

### v2

@@ Abstract
Although most current large multimodal models (LMMs) can already understand photos of natural scenes and portraits, their understanding of abstract images, e.g., charts, maps, or layouts, and visual reasoning capabilities remains quite rudimentary. They often struggle with simple daily tasks, such as reading time from a clock, understanding a flowchart, or planning a route using a road map. In light of this, we design a multi-modal self-instruct, utilizing large language models and their code capabilities to synthesize massive abstract images and visual reasoning instructions across daily scenarios. Our strategy effortlessly creates a multimodal benchmark with 11,193 instructions for eight visual scenarios: charts, tables, simulated maps, dashboards, flowcharts, relation graphs, floor plans, and visual puzzles. This benchmark, constructed with simple lines and geometric elements, exposes the shortcomings of most advanced LMMs like Claude-3.5-Sonnet and GPT-4o in abstract image understanding, spatial relations reasoning, and visual element induction. Besides, to verify the quality of our synthetic data, we fine-tune an LMM using 62,476 synthetic chart, table and road map instructions. The results demonstrate improved chart understanding and map navigation performance, and also demonstrate potential benefits for other visual reasoning tasks. Our code is available at: https://github.com/zwq2018/Multi-modal-Self-instruct.
@@
Wenqi Zhang1, Zhenglin Cheng1, Yuanyu He1, Mengna Wang2, Yongliang Shen1 Zeqi Tan1, Guiyang Hou1, Mingqian He1, Yanna Ma3, Weiming Lu1, Yueting Zhuang1 1College of Computer Science and Technology, Zhejiang University 2Institute of Software, Chinese Academy of Sciences 3University of Shanghai for Science and Technology {zhangwenqi, luwm}@zju.edu.cn Project Page: https://multi-modal-self-instruct.github.io
@@ 1 Introduction
To fill in the gap, we drew inspiration from synthetic data (Wang et al., 2022b; Liu et al., 2024c; Han et al., 2023; Du et al., 2023), which is widely used to supplement the insufficiency of instruction-following data. For instance, distilling high-quality dialogue data from a strong LLM (Wang et al., 2022b; Xu et al., 2023a; Yu et al., 2023a; Chen et al., 2023a; Zhao et al., 2023), or using external tools to refine the quality of synthetic data (Wei et al., 2023; Lee et al., 2024). However, synthesizing image-text data for LMM is not easy, as current LLMs can not directly generate images. An intuitive approach is to combine LLMs with a text-to-image model for producing <image, question, answer> (Li et al., 2023c; Wu et al., 2023b), but most text-to-image models fail to finely control the details of the image, potentially leading to a misalignment between image and text.
@@ 3.3 Implementation Details
To ensure the quality of the synthesized data, we filtered the data at three levels: code feasibility, image aesthetics, and answer accuracy. I. If the generated code fails to run, we prompt the LLM to self-reflect based on the error feedback from the compiler. If the LLM still cannot produce valid code after three retries, we discard that visual idea. II. For each synthesized image, we employed Llava-1.5 (Liu et al., 2024a) to check the image aesthetics, including whether visual elements within the image interfere with each other, the reasonableness of the layout, and the legibility of any text. These rules allowed us to filter out aesthetically unpleasing images. III. To ensure answer accuracy, we adopted the self-consistency (Wang et al., 2022a) for answer generation: instructing the LLM to generate multiple responses based on the idea, code, and question, and then selecting the final answer through a voting process. IV. Additionally, we randomly selected 10% of the question-answer pairs for human verification. The results confirmed that the quality of our dataset is assured.
@@ Appendix A Experiments Details
Appendix B Additional Experiment Results
As discussed in Section 4.2, we evaluate the performance of many LMMs, Llava-our-62k and humans using our benchmark. All results are shown in Table A1. Besides, as shown in Table B2, we also calculated the Rough-L score for the caption sub-task in the chart and table.
@@ Appendix B Additional Experiment Results
Table B2: For the chart and table tasks, we also calculated the captioning results.
Appendix C Related Work
C.1 Multi-modal LLMs
@@
C.2 Benchmark For Multimodal Model

### v4

@@ Abstract
Although most current large multimodal models (LMMs) can already understand photos of natural scenes and portraits, their understanding of abstract images, e.g., charts, maps, or layouts, and visual reasoning capabilities remains quite rudimentary. They often struggle with simple daily tasks, such as reading time from a clock, understanding a flowchart, or planning a route using a road map. In light of this, we design a multi-modal self-instruct pipeline, utilizing large language models and their code capabilities to synthesize massive abstract images and visual reasoning instructions across daily scenarios. Our strategy effortlessly creates a multimodal benchmark with 11,193 instructions for eight visual scenarios: charts, tables, simulated maps, dashboards, flowcharts, relation graphs, floor plans, and visual puzzles. This benchmark, constructed with simple lines and geometric elements, exposes the shortcomings of most advanced LMMs like Claude-3.5-Sonnet and GPT-4o in abstract image understanding, spatial relations reasoning, and visual element induction. Besides, to verify the quality of our synthetic data, we fine-tune an LMM using 62,476 synthetic chart, table and road map instructions. The results demonstrate improved chart understanding and map navigation performance, and also demonstrate potential benefits for other visual reasoning tasks. Our code is available at: https://github.com/zwq2018/Multi-modal-Self-instruct.
@@
Wenqi Zhang1,∗, Zhenglin Cheng1,∗, Yuanyu He1, Mengna Wang2, Yongliang Shen1 Zeqi Tan1, Guiyang Hou1, Mingqian He1, Yanna Ma3, Weiming Lu1, Yueting Zhuang1 1College of Computer Science and Technology, Zhejiang University 2Institute of Software, Chinese Academy of Sciences 3University of Shanghai for Science and Technology {zhangwenqi, luwm}@zju.edu.cn Project Page: https://multi-modal-self-instruct.github.io
11footnotetext: The first two authors have equal contributions.
@@ 1 Introduction
To fill in the gap, we drew inspiration from synthetic data (Wang et al., 2022b; Liu et al., 2024c; Han et al., 2023; Du et al., 2023), which is widely used to supplement the insufficiency of instruction-following data. For instance, distilling high-quality dialogue data from a strong LLM (Wang et al., 2022b; Xu et al., 2023a; Yu et al., 2023a; Chen et al., 2023a; Zhao et al., 2023), or using external tools to refine the quality of synthetic data (Wei et al., 2023; Lee et al., 2024). However, synthesizing image-text data for LMM is not easy, as current LLMs can not directly generate images. An intuitive approach is to combine LLMs with a text-to-image model for producing <image, question, answer> (Li et al., 2023c; Wu et al., 2023b), but most text-to-image models fail to finely control the details of the image (Betker et al., 2023; Esser et al., 2024), potentially leading to a misalignment between image and text.
@@ 3.3 Implementation Details
Dataset Diversity.
Firstly, in the data synthesis process, we control the generated topic of the image with many pre-defined keywords. For example, before synthesizing the chart, we designed a keyword library (e.g., GDP, energy, and employment rate) that includes various keywords from different domains covering economics, technology, and society. This strategy can control the generated content and avoid deviations. Similarly, during the image and question synthesis process, we use few-shot examples and templates to control the types of questions and images generated. For example, we generate five types of charts (bar, table, line, pie, composite) for the chart task, and also 5 types of questions (perception, extraction, math, caption, OCR). We also generate the difficulty levels of synthesized maps. The quantity for each category can be predefined in advance.
@@
To ensure the quality of the synthesized data, we filtered the data at three levels: code feasibility, image aesthetics, and answer accuracy. I. If the generated code fails to run, we prompt the LLM to self-reflect based on the error feedback from the compiler. If the LLM still cannot produce valid code after three retries, we discard that visual idea. II. For each synthesized image, we employed Llava-1.5 (Liu et al., 2024a) to check the image aesthetics, including whether visual elements within the image interfere with each other, the reasonableness of the layout, and the legibility of any text. These rules allowed us to filter out aesthetically unpleasing images. III. To ensure answer accuracy, we adopted the self-consistency (Wang et al., 2022a) for answer generation: instructing the LLM to generate multiple responses based on the idea, code, and question, and then selecting the final answer through a voting process.
Human Evaluation
we also conduct a manual evaluation of the dataset. First, we randomly sampled 10% of the <question, answer> pairs from our benchmarks and invited 4 graduate students in the computer science field for manual evaluation. For each sample, we designed four evaluation criteria: Image Aesthetics, Question Rationality, Answer Accuracy, and Image-Instruction Relevance. The criteria for Image Aesthetics and Answer Accuracy are scored from 1 to 5 (5 being the highest), while Question Rationality and Image-Instruction Relevance are divided into three levels 1, 3, 5. The scoring criteria for each dimension and the final results of the human evaluation are shown in Appendix B and Table B2.
@@ 4.4 Generalized to Untrained Tasks
4.5 How to improve the abstract image comprehension capabilities of LMMs?
More than just for instruction fine-tuning, our experiments have revealed that the abstract image comprehension capabilities of Large LMMs can be enhanced through various approaches:
Designing More Versatile Visual Encoders: First, we observe that current LMMs have weak visual representation abilities for abstract images, which may be caused by the current visual encoders. Most of them use clip-based encoders, which emphasize semantic features while neglecting purely visual features. We plan to explore adding another visual encoder, such as DinoV2, to Llava to improve its understanding of abstract images.
Increasing Image Resolution: Then we also observed that most current LMMs resize the original image to a resolution of 336x336 when training, as it reduces the number of visual tokens without losing much semantic information. However, for these abstract images composed of lines and geometric shapes, lowering the resolution results in the loss of a significant amount of geometric features, thereby affecting the ability of abstract visual perception. Increasing the image resolution and training LMMs from scratch may be a good solution.
Incorporate into Pre-training: Next, we will incorporate abstract image data during the pre-training stage, not just during the SFT stage. This is because, while benchmarking current LMMs, we found that their weakness lies in the perception of abstract images rather than in their instruction-following abilities.
Investigate Relationships between Abstract Image Tasks: Lastly, we will investigate the relationships between different abstract image tasks, quantitatively analyzing their mutual influences and their impact on LMM’s abilities such as abstract image perception, spatial reasoning, and visual-symbol Induction. These fine-grained studies will guide us in designing more useful abstract image tasks using our pipeline.
We believe that focusing on the aforementioned four aspects can effectively enhance the abstract image comprehension capabilities of LMMs without compromising their general capabilities. We treat this as a direction for future work.
@@ Appendix A Experiments Details
Appendix B Human Evaluation
As discussed in the paper, we design four evaluation metrics to manually assess the quality of the benchmark: Image Aesthetics, Question Rationality, Answer Accuracy, and Image–Instruction Relevance. The specific criteria are as follows:
  • Image Aesthetics: Are the colors appropriate, are the details clearly visible, is the spatial layout reasonable, and are there any obstructions between objects?
  • Question Rationality: Is the question reasonable? Is the wording clear and unambiguous? Is the difficulty level moderate, neither too simple nor too difficult?
  • Answer Accuracy: Is the answer accurate? Is the rationale logical? Are the answer and rationale consistent with each other?
  • Image–Instruction Relevance: Is the answer related to the image? Can the question be answered without looking at the image?
We evaluated the benchmark both before and after filtering, with the results presented in Table B2. These results indicate that the quality of our benchmark has significantly improved post-filtering, particularly in terms of Image Aesthetics and Answer Accuracy.
Image Aesthetics Question Rationality Answer Accuracy Image-Instruction Relevance
Before Filtering 2.4 3.9 3.5 4.5
After Filtering 4.0 4.1 4.3 4.4
Table B2: The results of the human evaluation.
Appendix C Additional Experiment Results
As discussed in Section 4.2, we evaluate the performance of many LMMs, Llava-our-62k and humans using our benchmark. All results are shown in Table A1. Besides, as shown in Table C3, we also calculated the Rough-L score for the caption sub-task in the chart and table.
@@ Appendix C Additional Experiment Results
Table C3: For the chart and table tasks, we also calculated the captioning results.
Appendix D Related Work
D.1 Multi-modal LLMs
@@
D.2 Benchmark For Multimodal Model